Subject: 3.5-m Users Committee minutes 3/31/08

From: Bruce Gillespie

Submitted: Thu, 3 Apr 2008 17:25:50 -0600

Message number: 1097 (previous: 1096, next: 1098 up: Index)

               APO 3.5-m Users Committee Phonecon, 3/31/08


Attending:  Suzanne Hawley, Bruce Gillespie, Scott Anderson, John  
Bally, Mark Klaene, Al Harper, Jon Holtzman, Jon Fulbright, Michael  
Strauss, John Wilson and Mike Skrutskie (both sitting in for Remy)

Absent: Remy Indebetouw, Russet McMillan

Minutes taken by Bruce Gillespie

**********************************


User feedback, comments from institutional representatives:

Princeton (Strauss) - Michael said the PU users were happy, and he had  
not heard of any problems or issues.

UVa (Indebetouw) - John Wilson (sitting in for Remy) had nothing new  
to report on user feedback or issues.

UW (Anderson) - Scott had nothing new to report on user feedback or  
issues.

CU (Bally) - John mentioned that CU has two Ph.D. theses finishing  
this semester, based on observations taken at APO.

JHU (Fulbright) - Jon had nothing new to report on user feedback or  
issues.

NMSU (Holtzman) - Jon had nothing new to report on user feedback or  
issues.

UChicago (Harper) - Al had nothing new to report on user feedback or  
issues.

**********************************


Discussion of telescope/instruments report - Klaene:

The report is given below.  Mark highlighted that the usual spring  
winds are kicking up dust, but the dust counts are not often reaching  
telescope closure levels (yet) as seen in previous years.  There has  
been little precipitation in the past couple of months, and the forest  
is extremely dry.  The Forest Service has just issued Level II (out of  
four) fire restrictions.  As precautions, the dead grass surrounding  
the site has been mowed, firelines have been cut, and the outside  
water sprinkler system has been set up.

The most significant recent issue with the telescope was the discovery  
and repair of a stuck secondary actuator, which seems to have been  
caused by a large ambient temperature swing.  All the instruments are  
in good shape, and TripleSpec was delivered and started its  
commissioning.  CorMASS was decommissioned and is being stored in the  
lower level of the telescope.  The tertiary support rebuild project  
held its critical design review, and its implementation appears to be  
on budget and schedule for completion during the summer shutdown.  The  
NA2 instrument rotator upgrade (both mechanical and controller) was  
completed and is now in service.  Detailed planning for the summer  
shutdown is underway.  Congratulations are in order for the APO  
technical and observing staff who contributed to all of these efforts  
during the past months, and to the TripleSpec development team.


                  *************************


3.5-m Telescope, Instruments, and CIF Projects Highlights, 2/021/08  
through 3/23/08

                      Mark Klaene & Bruce Gillespie


0) Overview

Operations continued largely routinely.  We are approaching the  
beginning of spring weather with the normal high winds and elevated  
dust levels, which will be monitored closely and may affect opening  
and closing conditions over the next few months.  The fire danger  
level in the surrounding forest is now very high and expected to reach  
the extreme level in upcoming weeks.  Planning and work is already  
being done to enhance the protection of the observatory and staff in  
the event of a nearby forest fire.  The SDSS-III construction for the  
expansion of the Support Building has been completed.

1) Telescope

Due to a stuck secondary axial actuator (thought to be caused by a  
significant ambient temperature swing), the telescope was out of  
collimation for a few days. The axial actuator problem was corrected  
quickly, however the transverse actuators while reporting good values,  
were also apparently mis-positioned.  Due to weather conditions at the  
time and confusion with the slit viewer being used to check  
collimation (which has its own intrinsic aberrations) the problem took  
a few days to find and correct.

2) Instruments

DIS, SPIcam, and the Echelle operated smoothly throughout the period.   
The new fanout board for NIC-FPS has been received at CU and we are  
waiting for them to schedule a date to warm the instrument, install  
the fanout board, and remove the etalon for testing back at CU.  We  
have proposed to do this work during time in April when NIC-FPS is not  
scheduled for observing.  With the arrival of TripleSpec, CorMASS has  
been decommissioned and is no longer available for observations.  The  
visiting instruments (GFP and APOLLO) were all operational.   
TripleSpec arrived at the observatory after a productive series of lab  
tests and cool-downs.  Initial on-sky commissioning started ahead of  
schedule because of the rotator upgrade work and ended with a  
successful three nights of commissioning.  Minor modifications will be  
required to make the instrument easier to mount and dismount, and TUI  
integration still needs to be completed along with improvements to  
guiding.

3) CIF projects

We are progressing on the implementation of Agile field rotation and  
its migration to the TR2 mirror cell port.  Design drawings have gone  
to the shop, and parts have been ordered or received.  The Critical  
Design Review for the Tertiary support/actuator upgrade was held on  
March 6 and 7th. The project cost estimates are within the planned  
budget, and the work is on schedule for a planned August  
installation.  The project was given a green light to proceed with  
only minor recommendations.  The new direct-drive rotator motor and  
axis controller were installed just prior to TripleSpec commissioning  
run.  Preliminary tests indicate that all objectives of the rotator  
drive upgrade have been met.  Further testing, analysis and cleanup of  
hardware and software will continue over the coming weeks.

4) Summer Shutdown

Planning is underway for the upcoming engineering shutdowns.  The  
initial take is looking like two 3-day shutdowns (in June and July),  
and a 2- to 3-week shutdown in mid-August.  All the shutdowns will be  
centered around bright time.

**********************************


TripleSpec commissioning - Skrutskie, Wilson, Hawley:

Suzanne was pleased to report that the delivery, installation,  
integration, and initial commissioning of TripleSpec has gone  
exceedingly well.  The main part that needs to be finished is the TUI  
interface software, after which we hope to release the instrument to  
our general user community.  This is likely to happen by May so that  
the instrument can be available for proposals for scheduled use in  
Q3.  We encourage prospective TripleSpec users to talk to Mike  
Skrutskie and John Wilson at UVa about using the instrument.  A  
preliminary user manual will be posted, and data reduction tools will  
soon be available.

Regarding instrument performance, Mike Skrutskie said that TripleSpec  
shows very low read or dark noise, and produces a decent S/N spectrum  
on a 15th magnitude target in a few minutes exposure.  Sky noise is  
the limit, as expected.  We intend to soon try some tests going  
deeper, to ~18th magnitude.

Suzanne asked that User Committee members inform their users of the  
upcoming availability of TripleSpec, and to encourage them to propose  
test targets to Mike and John for possible engineering test  
observations (shared-risk, non-proprietary) during April and May.  We  
also plan to announce that some of the TripleSpec commissioning data  
already taken will be made available to the ARC community so that  
people can learn more about the capabilities of the instrument.  As  
far as we know, only APO and Palomar have mid-sized telescopes that  
can take single-exposure, full-wavelength coverage NIR spectra.   
Again, our congratulations and appreciation go out to the UVa  
instrument developers and the APO staff for their hard work in  
reaching this outstanding accomplishment.

**********************************


NIC-FPS status - Bally:

John said that the new, low-noise, fan-out board for NIC-FPS has been  
fabricated.  The schedule and arrangements to install the board at the  
site are being worked out, hopefully in May.  At the same time as the  
fan-out board is replaced, the faulty Fabry-Perot etalon will be  
removed and returned to CU for analysis and repair.

**********************************


Instrument rotator upgrade - Klaene:

The upgrade to the NA2 instrument rotator was discussed in detail  
above in the Telescope/Instrument report.  So far, it appears to be  
working very well (with higher torque and better motion control than  
the old rotator system); we ask that if any of our users see quirks in  
its behavior during their observing runs to please notify the APO  
staff.  Bruce said that although this was largely an APO effort (kudos  
to Nick McDonald and Fritz Stauffer), we should also acknowledge and  
thank UVa for their significant support for this upgrade, which was in  
part needed to accommodate Triplespec.

**********************************


Tertiary mount CDR, and autofocus - Klaene:

Although also already discussed above, Mark added that the rebuild of  
the tertiary mount will be fairly straightforward (compared to the  
past rebuild of the top end and secondary mounts).  When implemented  
this summer, we should see improvements in stiffness, mirror-position  
stability, maintainability, and safety to the glass.

On the autofocus project, substantial progress is expected in the  
coming months, which will be reported on in future meetings.

**********************************


Curation of NIC-FPS raw data - Hawley:

Suzanne reminded the committee to inform their users that we are about  
to implement the policy of deleting all raw NIC-FPS data older than 90  
days. This was discussed in detail at last month's meeting, and in  
early April we will purge all NIC-FPS raw data (but not the processed  
images) stored at APO that were taken before January '08.  A separate  
announcement to this effect will be posted to apo35-general.

**********************************


ACTION ITEMS:

[open from previous months]:

ACTION:  John Bally will write a section with figures on Fowler  
sampling for inclusion in the NIC-FPS user documentation.

STATUS:  John will check to see if the Fowler sampling documentation  
has been posted and is adequate.


[new actions from this month]:

none

**********************************


Next meeting:

The next Users Committee phonecon will be on Monday, 28 April, at 8:30  
AM Pacific Time.  The agenda and other materials will be sent to the  
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO
APO  This is message 1097 in the apo35-general archive. You can find
APO  the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO  To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO  To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO